solve the system
x^2+y^2=4
4x^2+9y^2=36

Answer by oscargut(900)

x^2+y^2=4
4x^2+9y^2=36
using 1st eq y^2=4-x^2
using the 2nd one
4x^2+9(4-x^2)=36
-5x^2+36=36
-5x^2=0
then x=0
then
y^2=4
and y=2 or y=-2
Answer: solutions are: x=0,y=2 and x=0,y=-2